### v1

# Large Language Models in Qualitative Research: Can We Do the Data Justice?

Abstract.
Qualitative researchers use tools to collect, sort, and analyze their data. Should qualitative researchers use large language models (LLMs) as part of their practice? LLMs could augment qualitative research, but it is unclear if their use is appropriate, ethical, or aligned with qualitative researchers’ goals and values. We interviewed twenty qualitative researchers to investigate these tensions. Many participants see LLMs as promising interlocutors with attractive use cases across the stages of research, but wrestle with their performance and appropriateness. Participants surface concerns regarding the use of LLMs while protecting participant interests, and call attention to an urgent lack of norms and tooling to guide the ethical use of LLMs in research. Given the importance of qualitative methods to human-computer interaction, we use the tensions surfaced by our participants to outline guidelines for researchers considering using LLMs in qualitative research and design principles for LLM-assisted qualitative data analysis tools.
Qualitative methods, large language models
journalyear: 2024copyright: rightsretainedconference: Extended Abstracts of the CHI Conference on Human Factors in Computing Systems; April 26-May 1, 2025; Yokohama, Japanbooktitle: Extended Abstracts of the CHI Conference on Human Factors in Computing Systems (CHI EA ’25), April 26-May 1, 2025, Yokohama, Japandoi: 10.1145/3613905.3636301isbn: 979-8-4007-0331-7/24/05ccs: Human-centered computing HCI design and evaluation methodsccs: Human-centered computing HCI design and evaluation methods
1. Introduction
The qualitative research process is traditionally grounded in deep engagement with participants and their data through an iterative sensemaking process. Today, Large Language Models (LLMs), particularly those accessed through chat-based interfaces like ChatGPT  (OpenAI, 2024), are popular tools in research-related tasks, including in the social sciences (Grossmann et al., 2023). While a growing number of dedicated tools for qualitative research use LLMs to streamline segments of the qualitative research process, like scaffolding thematic analysis or automating qualitative coding  (Gebreegziabher et al., 2023; Gao et al., 2024; Choksi et al., 2024), direct interactions with ChatGPT and similar interfaces may be increasingly used by qualitative researchers as well, a topic which has not received thorough investigation to this point.
LLMs distinguish themselves from prior tools for qualitative research for several reasons. First, LLMs are not dedicated qualitative analysis tools. Unlike most dedicated software platforms, users interact with them in natural language, leading to highly fluent interactions. LLMs go beyond simple document-organization or analysis functions in classic tools, and appear to facilitate highly flexible interpretations of documents, such as the ability to produce summaries or annotations of documents, behaviors which mimic aspects of the human sensemaking process core to qualitative research. Second, LLM use is currently dominated by one product, ChatGPT, that has rapidly entered general use by the public (Porter, [n. d.]), including qualitative researchers. Third, the dominant user interface metaphor (Neale and Carroll, 1997) of ChatGPT and its competitors is that of a jack-of-all-trades assistant and/or collaborator. This pattern is simultaneously intuitive thanks to the conversational interface, yet completely unguided. With wide-ranging and open-ended interactions with LLMs increasingly permeating everyday life, qualitative researchers are left to wonder whether or how these tools might apply to their research process or outputs. While some qualitative researchers support integrating LLMs into aspects of qualitative research, others, as  Soden et al. (2024) highlight, will be “concerned that increasingly all qualitative research, including that which draws on interpretive traditions of research design, is being evaluated from the perspective of positivism.”
There is a critical and immediate need to understand the degree to which LLMs are shaping qualitative research processes. On the one hand, many researchers may benefit from the inclusion of LLM-based features in qualitative data analysis (QDA) systems, as has been explored in HCI literature (Gao et al., 2024; Jiang et al., 2021; Overney et al., 2024). However, the recent explosion of work using LLMs to annotate text in ways previously difficult for advanced models (Ziems et al., 2024; Bail, 2024) may risk the excitement of communities of scholars and tool builders without full consideration for potential tradeoffs to the sensemaking process. In particular, readily available tools that seek to automate the qualitative sensemaking process could decrease the amount of time researchers spend with their participant data, a potential risk to qualitative researchers’ depth of understanding.
Furthermore, the speed of LLM development has outpaced guidance on their ethical use (Baronchelli, 2024). Most available policies focus on the danger of uploading sensitive information to cloud services (Ollion et al., 2024) or text generation in publications (Hosseini and Horbach, 2023). The HCI community specifically is contending with the need for developed policies that tackle the challenge of how to use AI ethically in research (Shen et al., 2023; Aubin Le Quéré et al., 2024). While guidelines are beginning to emerge, there are no guidelines specific to qualitative work today.
This work critically reflects on the role of LLMs in qualitative research and offers guidance to researchers considering using LLMs. Through our work, we answer the following research questions:
  • In what ways are LLMs being adopted into qualitative research processes today, and how might they be incorporated in the future?
  • What do qualitative researchers say they stand to lose or gain from adopting LLMs into their research processes, and what tensions might arise?
  • What are potential models for the ethical incorporation of LLMs into the qualitative research process?
Qualitative research is a broad and rich tradition with scholars from numerous disciplines and ways of thinking. This paper focuses on the subset of researchers, including mixed-method researchers, who are considering using LLMs in their qualitative research processes. Those researchers should be clear-eyed and cautious about what is gained and lost when LLMs are used in data collection, analysis, or writing.
This work surfaces both the curiosities and concerns of qualitative and mixed-method researchers across disciplines who are wrestling with what the introduction of LLMs means for their qualitative research process. Through twenty semi-structured interviews with researchers across academic communities, we sought to understand any current use, motivations for use, perceptions of usefulness, and concerns about the use of LLMs in qualitative research. Our findings show that qualitative researchers can envision leveraging LLMs for their work to help with a range of uses including generating interview materials and analyzing research artifacts. Specifically, the interactivity of ChatGPT has spurred researchers to think creatively about what chat-based interfaces can bring to their work. Simultaneously, researchers harbor fundamental concerns about ethics, unequal adoption of new technologies, model bias, and performance, leading some to abstain from the use of LLMs to conduct qualitative research and others to experiment despite concerns to participant privacy. New tools and practices are rapidly being adopted, but scholarly communities are not equipped for conducting ethical and effective research with them. As such, we are at a critical junction that requires urgent evaluation and recalibration of both our methodological approaches and thinking on research practices..
As a step forward, we outline guidelines for future qualitative researchers interested in using LLMs, and we present design opportunities and principles for LLM-based qualitative data analysis tools.
2. Related Work
Qualitative research is a broad term with different definitions and applications across disciplines (Hammersley, 2013). In the late 1960s, the method first emerged in retaliation to quantitativeand positivist approaches to science characterized by hypothesis-testing and objectivity (Hammersley and Richardson, 1996). HCI scholars often turn to qualitative methods when studying populations or contexts that require an in-depth exploration of experiences, behaviors, and interactions (Randazzo and Ammari, 2023; Randazzo et al., 2023; Rankin and Han, 2019; Freeman and Wohn, 2020; Bardzell and Bardzell, 2011b, a).
In contrast to quantitative methods that prioritize numerical data and generalizability, qualitative research aims to uncover how participants perceive, understand, and make meaning of their experiences (Jaccard and Jacoby, 2020). Qualitative researchers often utilize diverse techniques, such as interviews, observations, and narratives (Alvarez et al., 2023; Nolan-Cody and Scharp, 2024) to capture the complexity of lived experiences (Braun and Clarke, 2013). The epistemological foundation of qualitative research is social constructionism (Berger and Luckmann, 1966), in which meaning and knowledge is subjective and constructed in and through communication (Weinberg, 2009), enabling researchers to capture the social contexts that shape participants’ lives (Tracy, 2024). Through deep engagement with participants’ worlds, qualitative scholars seek to describe and interpret phenomena in a way that reflects participants’ understanding of their experiences (Jaccard and Jacoby, 2020), while also contributing to theory development through inductive analysis (Scharp, 2021).
As a field, HCI hosts strong qualitative approaches but also mixed-method research, which has sometimes led to tensions in how contributions should be evaluated in the field. In particular, a stigma surrounding the perceived rigor of qualitative research (Charmaz and Thornberg, 2021) has sparked debates over how best to evaluate such methods (Kvale, 1995; McDonald et al., 2019; Rolfe, 2006). As a result, cting qualitative research can involve a complex and sometimes creative negotiation with the norms and evaluative criteria of one’s research communities (Denzin et al., 2024). Within the HCI community, there have been concerns that increasingly quantitative evaluation techniques, such as providing participant demographics, are being unduly required of qualitative scholars (Soden et al., 2024). LLMs may lend the impression that qualitative inquiry can be automated, and their integration into Qualitative Data Analysis (QDA) software may increasingly impose positivist approaches that conflict with interpretive traditions.
Approaches to analyzing qualitative data (e.g., grounded theory, thematic analysis) were originally designed for manual, pen-and-paper applications (Deterding and Waters, 2021). However, the widespread adoption of software tools for qualitative data analysis marked a significant and, at times, contentious shift in the field (Gilbert et al., 2014). Early critiques of qualitative software centered around concerns that it created too much distance from the data, compromised the depth of analysis (Blismas and Dainty, 2003), or could be misused by researchers who lacked a reflective understanding of them (Gilbert et al., 2014).
Advancements in natural language processing and machine learning over the last decades have widened possibilities for data analysis. LLMs are quickly being adapted for text annotation in increasingly interpretive contexts (Ziems et al., 2024; Bail, 2024). In HCI, tools that leverage these advancements have been developed to assist qualitative researchers since before the most recent LLM era. Past published work has surfaced qualitative HCI researchers’ imagined uses of AI in analysis, and found researchers were open to using AI in data exploration and annotation with stipulations regarding researcher agency (Feuston and Brubaker, 2021) or to “support serendipity” in the analysis process (Jiang et al., 2021), while also surfacing concerns about potential tradeoffs for qualitative research these tools could present. Past work has also considered the tensions and potential compatibilities between computational tools and interpretive work (Baumer et al., 2017; Nelson, 2020). More recent tools have built on this work, leveraging a new generation of LLMs to create tools for qualitative researchers that bear researcher goals in mind, like agency over the coding process (Overney et al., 2024) or convergent discussion on coding among collaborators (Gao et al., 2024). Each contains insights from users that give insight into the promises and pitfalls of LLMs in qualitative research, but evaluations focus on users’ perceptions of particular tools. Given the sudden, widespread ubiquity of LLMs through task-agnostic interfaces like ChatGPT, experiences with AI tools have become more common for all kinds of users, including those not in computing. The sudden uptake of LLMs in society has also changed common perceptions about which uses cases may be possible in the future, motivating a fresh examination of current uses and attitudes towards LLMs in qualitative research that also includes researchers outside of HCI.
3. Methods
To gain a deeper understanding of the experiences qualitative researchers are having with generative AI tools, we conducted a series of twenty semi-structured interviews in June-August 2024. We analyzed the interviews using inductive thematic analysis, inspired by grounded theory  (Braun and Clarke, 2006), in which concepts emerge from the data (Jaccard and Jacoby, 2020). These interview procedures were approved by the Institutional Review Board at [anonymous institutions].
Participants were recruited mostly based on personal networks, past work, and through in-person events. Members of the research team also participated in in-person events, where they were able to conduct interviews on-site. On some occasions, we encountered participants’ prior works online and emailed them to request an interview. The inclusion criterion for this interview was simple: we sought to speak with researchers who consider themselves qualitative scholars, including researchers who use qualitative methods as part of a mixed methods approach to their work. Since we sought to maximize variety of perspectives, we did not recruit participants based on their use or knowledge of LLMs, and included participants across academic domains. Since qualitative research perspectives are field-dependent, we considered it important to allow potential interviewees to self-define as qualitative researchers. These recruitment strategies resulted in twenty participant interviews, summarized in Table 1.
The interview protocol was designed to elicit perspectives on LLMs in qualitative research, with a focus on data collection and analysis processes, as they are more unique to qualitative research than other steps of the research process, which have been the focus of other work (Kapania et al., 2024). For this study, data collection was defined as any part of the research process that contributes towards how researchers obtain data from participants. LLMs may play an active role in data collection, e.g. through creating recruitment materials for different formats. Data analysis was defined as any part of the research process that relates to the analysis of gathered data. LLMs may play a role in data analysis, e.g. through applying qualitative codes. We opened the interview with questions about the researcher’s qualitative work, practices, and their general familiarity with large language models. We then asked questions about data collection and analysis, asking the researcher if they have or could imagine using LLMs for any portion of these research processes. If they said they had used LLMs for that purpose, we asked them about their motivations, experiences, and ethical considerations. If they had not used LLMs before for that purpose, we asked them why, and whether they had major concerns with others using LLMs for these purposes. We concluded our interviews by discussing the participants’ perceptions of LLMs in qualitative research, soliciting potential use cases and concerns for using LLMs in qualitative research. The full protocol is included in the Appendix.
We analyzed the data using inductive thematic analysis techniques inspired by grounded theory (Braun and Clarke, 2006). Once the interviews were recorded, three researchers open coded two interviews each, and each compiled a proposed codebook structure, which we consolidated through discussion. Two further interviews were fully coded with this first-round codebook, and we iterated once more on the codebook to yield the final codebook. We then divided the twenty interviews among coders, and met regularly to discuss the process. We used Dedoose for collaborative code application to interview transcripts. We primarily avoided using LLMs for this research, with the exception of finding specific works of related literature, which we then verified externally. To protect the identity of participants, any specific research studies or topics have been changed.
We take a moment to reflect on how our own positionality influenced this study. The lead authors of this study are mixed-method researchers in HCI and communication. In our own work, we have felt a desire and some pressure to incorporate LLMs into our research processes, while being unsure how to ethically navigate adopting this technology. This work is an effort to tackle these difficulties head-on and to shine a light on the value of qualitative work in HCI.
Participant ID Academic Domain Position Approach to Research
P1 HCI PhD Student Mixed-Method
P2 Communication PhD Student Mixed-Method
P3 Political Science Research project manager Primarily Qualitative
P4 HCI PhD Student Mixed-Method
P5 Social Work PhD Student Mixed-Method
P6 Communication PhD Student Primarily Qualitative
P7 Political Science PhD Student Mixed-Method
P8 HCI PhD Student Primarily Qualitative
P9 HCI Professor Mixed-Method
P10 HCI PhD Student Primarily Qualitative
P11 HCI PhD Student Primarily Qualitative
P12 HCI PhD Student Primarily Qualitative
P13 Sociology Research Scholar Mixed-Method
P14 Political Science PhD Student Mixed-Method
P15 Sociology Former Professor Primarily Qualitative
P16 Communication PhD Student Primarily Qualitative
P17 Sociology Post Doc Mixed-Method
P18 Political Science PhD Student Mixed-Method
P19 Anthropology PhD Student Primarily Qualitative
P20 Psychology PhD Student Mixed-Method
Table 1. Participant Summary Table. Participants self-reported their academic domains, position, and approach to research.
4. Findings
In the findings section, we dive into how the researchers we interviewed that adopted LLMs into their qualitative research processes, we identify the main risks and concerns they raised, and conclude by outlining the fundamental philosophical tensions that arise in considering how qualitative work may be combined with LLMs.
4.1. Current and Potential Uses of Qualitative Researchers Adopting LLMs
[…]
Current and Envisioned Use Cases for LLMs within the Qualitative Research Process. This figure highlights how participants had previously used and might imagine using LLMs moving forward in their research processes. To the left, the use cases under ”LLMs in Data Collection” are Creating/editing study materials and Adapting materials to audiences. In the middle, the use cases under LLMs in Data Analysis are Data Transformation & Retrieval, Data exploration, Generating qualitative codes & themes, and Applying qualitative codes. In the right, the use cases under ”LLMs in Other Research Tasks” are Ideating, Writing & editing, Finding related work, and Learning & Coding. Each use case is accompanied by a participant quote that supports and refutes the use case.
Figure 1. Current and Envisioned Use Cases for LLMs within the Qualitative Research Process. The figure highlights how participants had previously used and might imagine using LLMs moving forward in their research processes. We also highlight different perspectives and tensions for each use case, with participant quotes. Participants are only tagged under a use case if they view LLMs as being potentially beneficial to that end, and use cases are only included if mentioned by three or more participants.
We sought to understand how our participants had experimented, or not, with using LLMs as part of their qualitative research process. Of the researchers we interviewed, fourteen reported having actively explored the use of LLMs in data collection or analysis. The most common use cases mentioned, highlighted in Figure 1, were using LLMs to help generate recruitment materials, attempts to use LLMs to speed up qualitative coding, and using LLM to inspire or provide feedback on ideas. In this section, we foreground perceived potential uses of these methods, before we platform the perceptions and risks of their use in the next.
Unsurprisingly, ChatGPT was the most frequently mentioned LLM tool by participants. When we opened the conversation by asking participants about their experiences with “generative AI tools or large language models,” most would immediately talk about ChatGPT. Some researchers specified that there is a distinction; for example, P11 opened by saying “I use LLMs, specifically ChatGPT.” Others spoke more hesitantly, such as P13: “I’ve started using ChatGPT, I’m trying to be more familiar with it because it is the future, and you don’t want to be a dinosaur.”
Outside of ChatGPT, researchers noted that they used a suite of available tools for qualitative research, such as NVivo, Atlas.ti, or Dedoose, and AI sometimes underlies these tools. Within these tools, participants were sometimes unsure whether they were leveraging LLM-based functionality. For example, participants (P4, P8, P9, P13) wondered: “does Grammarly count?” Others mentioned explicitly that they engaged in AI features offered by these platforms, such as P7, who explained that through the tool Atlas.ti “people are getting suggestions […] in terms of coding certain paragraphs or sentences as in our text or document.” Participants did not necessarily feel agency over the intelligence embedded in these tools, as P1 explained: “Atlas sends me a thematic analysis that essentially aggregates stuff, I don’t really like it,” and P4 summarized AI features in these tools as “not by choice.” Overall, participants were used to relying on technology platforms to help with research, though not all appreciated the increased inclusion of smart features.
In contrast, some researchers mentioned centering a purposeful lack of advanced technology in their qualitative process. While all participants used their computers for analysis, they also talked about paring back their approach to remain close to the data. P6 explained that while they were “not quite as old school as the people who would cut out their coded texts and move it around, I basically do that in Word documents. […] I don’t even use a qualitative management software.” A similar sentiment was echoed by P2, who recounted “I didn’t like the restrictions and things the platform was putting on me, and I got really frustrated; I’m just familiar with Excel.” These attitudes highlight how, for some researchers, any software interferes with their process.
4.1.1. LLMs during Data Collection
Seven participants reported using LLMs to some degree to help with their recruitment or interview materials. Often, this took the form of researchers using a service like ChatGPT to edit or tweak their writing. Even those who were reluctant to use LLMs for research conceded that they could be helpful for writing, such as P10 who said “the only real reason that I use ChatGPT is to have it reword or rephrase,” and P4 similarly explained that “it will give me a really good sentence or switch a word.” This capability was particularly helpful for participants whose native language is not English; for example, P8 said they “feel more confident once I have passed my writing to the AI tools.” LLMs also assisted participants in tailoring recruitment materials to their subject population. One notable example was P9, a non-American researcher whose first language is not English, who told us they used LLMs to make messaging more culturally sensitive to research participants. They shared, “one thing I noticed that LLMs can do beautifully is make things American,” and therefore “an LLM’s general answer or general email to an American future participant will probably be better than my take on it.” P5 reported that LLMs helped them deploy recruitment materials in multiple languages, explaining that they “used it for the translation of interviews and consent forms and other required documents,” before having a native speaking collaborator double check the output. LLMs were also used by participants to help them create interview and recruitment materials from scratch. For example, P8 recalled “I was working with a deadline, and I was feeling stuck. I was not sure what kinds of questions to ask, but I did have a set of broad topics […], so I put those into ChatGPT and I was like, ‘Can you help me come up with this type of interview question?”’ While our participants had experimented with using these tools to help generate research materials, they did not all find it helpful: “the performance was really great,” P5 said, while P8 thought “I feel like I could have done better,” and P10 elaborated that outputs were “really broad, versus the empirical work I’ve done is very specific and granular.”
In terms of promising potential uses, researchers seemed to understand why one might use LLMs to help create recruitment materials, even when they did not use them themselves. Some experienced participants explained that they had their own past templates, but P10 added “I can imagine that maybe using something like this would be helpful for beginner researchers.” P20 expressed a similar feeling: “I would consider it in the future, if I’m drafting a new consent letters or things like that, but we had all of that from previous ones.” Nonetheless, some researchers pointed to specific reasons why they may not consider LLMs always to be appropriate for these use cases, such as when materials are being produced for at-risk or non-Western populations, which will be further discussed in the following sections
4.1.2. LLMs during Data Analysis
Another avenue that participants had explored, though often less successfully, was using an LLM to make process and make sense of their qualitative data. Many researchers we talked to did not attempt such analysis, since they struggled to understand whether they could share even anonymized participant data with third-party LLMs, an issue we explore in depth in the next section. Among those interested in LLM-facilitated analysis of qualitative data, we saw a few functions being performed with LLMs: transforming or retrieving data, data exploration, generating initial codes or themes, and applying qualitative codes.
Researchers wanted to use LLMs as assistants for retrieving, processing, transforming, or reformatting their data using natural language. P9 reported that they leverage LLMs to perform fuzzy searches on qualitative data: “With retrieval via an LLM, I can say something like, ‘Where in this document is an example of such and such?’ and just talk more conceptually or more abstractly, and the tool will find it.” Participants also saw LLMs as being potential useful to reformat data, for example P4 imagined that “it would be nice if there was like a large language model agent to grab all the qualitative codes [and put them] onto a Google Jamboard because I have to rely on copy pasting.”
On the analysis front, participants explained that they valued being able to use LLM-aided software to initially explore participant data. In P15’s words, the ability to easily analyze qualitative data with a sensemaking model was a “game changer,” since “there’s no friction between the person trying to make sense of a lot of data and the information.” One participant said they once used ChatGPT to help them make sense of the theory behind their qualitative data, explaining that they had interviewed people about their experiences with grief. They were “getting all this lovely, well-gelled data where people were repeatedly saying the same things, everything was perfectly categorized” but were “having a hard time naming those categories in a way that was as inclusive as it could be.” In this example, the participant asked ChatGPT to come up with a list of potential themes, based not on their participant data but on prior theories, and it gave them “this list of things, and I was like, hmm, this is really neatly packaged […] it’s giving me labels for the things I’m seeing that are also very intuitive.” These themes then served as a jumping off point for deeper analysis.
Experiences were more mixed for participants who directly tried to use ChatGPT or other LLMs to assign qualitative codes to data. The most common experience was perhaps that of P5, who said “I tried to use ChatGPT to get the very general theme and the results were not that great, so I decided to code it manually, but I tried it.” P9 engages in a process they refer to as “semi-coding,” where they provide ChatGPT a chunk of text they already had intuitions about, and wanted to see if the LLM agreed with their perspective. Ultimately, however, they felt that the codes were “uninteresting in a way, or at least they were first level codes.” This experience is also shared by P11, who explained that when they “asked it to assign some initial codes [on their dataset], oftentimes the response will be very high level and abstract, so it’s not trustworthy at all.” P17 disagreed, and had set up specific coding infrastructure that allowed them to identify whether an individual piece of social media data contains a rare occurrence of interest to their research. P17 was seeking to identify and code specific types of gender-based online harassment in Reddit comments and derived the most advanced annotation pipeline we saw:
“I come up with 16 or 18 thematic categories, and then use them to describe the system that I’m looking at. That’s qualitative work… I have… millions of comments, and most of those are garbage… So it’s a needle in a haystack, and computation allows you to find and organize all those needles so that you can do the qualitative work.”
In their workflow, P17 derived codes based on initial data, and then constructed an LLM pipeline to annotate each post, using the LLM to take a first cut at organizing the data. P17 expressed that this process does not preclude deep qualitative work, since they still do a manual layer of qualitative analysis.
Despite overall mixed current experiences stemming from perceptions of ChatGPT, and privacy concerns with today’s systems, many participants felt curious or excited by the prospect of using LLMs to automate qualitative coding. The biggest factor that motivated curiosity around the use of LLMs for annotation was the tedium and time investment required by manual coding. P10 said “I’m currently in the process of coding, and sometimes I’m kind of like, it’s taking a long time.” Others mentioned that automating coding could limit human error and help them optimize for inter-rater reliability. P13 coordinated a large project with multiple coders, and they described that “having a more thoughtful upfront conversation about that, and then using a tool to execute on that [code scheme], I think, would be way more accurate and better science than what we ended up doing.” Of those interested in using LLMs for coding, researchers had conflicting perspectives about whether LLMs would best be used for an initial, exploratory round of coding, or whether they should be used to apply manually developed codes. Some specifically identified that LLM-aided coding could help with the first round of coding, or when initially getting a grip on the data. P13 thought LLMs could “save you from that first tedious round of coding and then you could maybe be looking at the more nuanced codes faster over time.” Others described the desire to use LLMs as a way to “gut check” their coding. For example, P18 explained that they had considered if “me and another human are coding the data, and then ChatGPT is also coding the data, and we do an intercoder reliability check with all of us.” Nonetheless, some participants (P2, P4, P8, P16) struggled to see the automation of coding as aligned with perception of qualitative work, stating “there’s a part of qualitative research where it kind of feels like I’m in this for the manual work and that’s why I’m doing it” (P2).
4.1.3. LLMs in Other Research Tasks
Beyond data collection and analysis, participants also reported leveraging the interactivity of readily available LLMs, particularly ChatGPT. Participants felt that ChatGPT could help them ideate better and be more creative researchers; “I think LLMs are really great for bouncing ideas off of and thinking creatively,” P18 said, while P16 said they “will use it as like a basis or a starting point.” Participants also explained that they used ChatGPT to generate explanations for complicated concepts or coding implementations, like P7, who explained that they ask ChatGPT to help explain specific theories, stating “if I didn’t understand what kind of methods they used in a particular paper or […] I’m not able to make some connections or linkages, I just pose a question to ChatGPT.” In these instances, the social semblance of LLMs takes center stage, and our participants would sometimes refer to ChatGPT as more of a collaborator or peer than a tool. For example, P18 referred to ChatGPT as a “great brainstorming buddy,” P20 called it a “thought partner”, P9 defined it as “a rational person, a smart layman, if you will,” and P12 said that working with ChatGPT was “almost like working with another person.” For P11, this connection went further, as they found that interacting with ChatGPT while doing research helped stave off loneliness:
“As a researcher, we are always heads down and maybe we do not have much interaction with peer researchers who are human, o I feel like I really value this kind of interaction, somebody I can talk to. I feel like this research process is not boring, it’s engaging.”
However, some researchers were more critical of ChatGPT occupying these kinds of interpersonal roles. P5 expressed frustration that ChatGPT does not push back enough on their ideas to be a collaborator, since ChatGPT “never criticizes me, even though I asked: pretty, pretty please criticize me.” P4, meanwhile, was dismayed at the prospect of replacing any of their processes with ChatGPT, saying “I have a very supportive advisor and group of collaborators; they will give me great human feedback and I wouldn’t I even think about using language models for that purpose.”
4.1.4. Viewing LLMs and Generative AI through the Lens of ChatGPT
Overall, perceptions of LLMs were strongly colored by ChatGPT’s performance as the most popular interface through which to explore LLMs. For example, this conflation impacted people’s general perception of LLMs’ suitability for complex tasks. Some participants tried to use ChatGPT for data annotation, observed mediocre results (“vague” (P6), “too generic” (P5), “very general” (P13), or “uninteresting” (P9)), and thus reached the conclusion that LLMs could not help their use case. However, some participants looked beyond out-of-the-box ChatGPT. For example, P17 set up a customized labeling pipeline that required some engineering, and P11 used a personal version of ChatGPT to improve results. By the same token, participants’ ideas about what might now be possible for the next generation of data analysis tools were likely expanded by ChatGPT and imagined natural language interfaces for tasks like data transformation and retrieval without technical knowledge. Generally speaking, participants felt that LLMs were more advanced in areas where ChatGPT performs better, thanks to their interactions directly through its web interface – for example, more participants felt that LLMs were helpful “brainstorming buddies” (P20), than felt they were helpful annotators, which is a use case for which ChatGPT in the web interface was not designed. Similarly, people’s perceptions of the privacy implications of sharing participant data with LLMs more broadly were influenced by their knowledge of ChatGPT and OpenAI. P16 demonstrated this tension nicely when they advised, “Don’t put anything your participants have said into LLMs. I don’t know much about it, but I’m guessing it collects the data.” This is an example of how multiple participants reported that they believed it unethical to share participant data with a third-party tool like ChatGPT, but did not discuss the potential alternative of open-source models to address some concerns. This may be due to a combined lack of knowledge on the subject or the relative scarcity of practically implemented alternatives for users, particularly those without technical backgrounds.
4.2. Perceptions and Risks of LLM Use in Qualitative Research
While researchers identified potential use cases for LLMs in research, and expressed curiosity about new methods, most had not used LLMs extensively in research. During our interviews, participants shed light on barriers to adoption, including a lack of established norms and concerns about output quality and appropriateness.
4.2.1. Uneven adoption
Researchers reported logistical hurdles to adopting AI tools, which seemed related to their comfort with technology and inequality around funding. For example, P13 expressed that the prevalent feeling among their team when faced with a new tool is: “We’re going to have to learn this, it’s gonna take forever, and we’re never going to get around to doing the actual project.” P2 and P5 recounted that cost influenced their choice of models, P2 stating “since the Claude model is quite expensive, I didn’t use it, and OpenAI models are much cheaper, so I could test it out.” Researchers also may not be certain how to find tools that meet their needs; P20 described how “you don’t know what’s out there, sometimes you don’t even know what something’s called.” Overall, logistical issues led researchers to be more reliant on accessible interfaces like ChatGPT, even when these are not optimized for their use case, an issue exacerbated for non-technical researchers.
4.2.2. Uncertainty about norms, guidelines, and expectations
Paricipants shared that university or publication venue policies are shaping their perception of what is acceptable (P4, P11, P14), but many guidelines fall short. For example, participants referenced that university policies surrounding LLMs to produce writing informed their decisions about using LLMs for this use case, but also fuel uncertainty about using LLMs for other purposes. P14 reflected, “ I guess, as students, … it’s kind of concerning that we might be penalized for using ChatGPT, so we’re probably very, very careful” and “I don’t want this to backfire and cause any problems.” P2 elaborated: “I get nervous… whenever it’s not me, when I feel like it’s not me, is someone gonna say that I’m plagiarizing something?” Sixteen participants in our study were graduate students, who occupy the unique role of collecting and analyzing study data directly, while feeling little power over the direct creation of guidelines. Some researchers mention efforts to seek additional clarification, such as P13 who reached out to legal experts in the university to clarify whether they are allowed to “use AI” when analyzing their data.
Participants also raise concerns that LLMs may introduce IP and originality issues that make them cautious about their use in research (P11, P14, P18, P20). For example, P11 shared, “I feel like the IP is kind of mixed up in the data collection, because ChatGPT does give me some creative answers, but then sometimes I feel like I just take it for granted, and I probably did not think much about whether this creative answer has its own intellectual property or not.” Continuing, they asked, ”Is it bad to ask it to do stuff, and then kind of look it over, does that still count as your material? I think we’re getting into those waters now.” Others note that as LLMs are increasingly integrated into workflows, policies that require disclosure, like the ACM policy, may be hard to enforce in practice. P18 wrestled with this concept, too: “At some point it may be somewhat impossible to truly distinguish what a human contributed or thought about, from what an AI that they were working with contributed.”
4.2.3. Navigating Participant Privacy Implications
Our participants were most concerned with how LLM use impacts data security, consent, and obligations to participants in qualitative research.
Participants expressed a general lack of understanding and distrust for data protections awarded to participant data sent to proprietary models. For example, P7 expresses their thought process: “How will the companies use that data […] how is their data management protocol? Because the data that we share with LLMs, we don’t know how that is managed at the back end.”
Researchers also wondered whether classic models of participant consent that cover sharing data with third party proprietary models like ChatGPT are doing enough to protect participants, apply, like P8: “There’s a bigger question of … primacy of what is going to be giving consent and whatnot.” Some participants held strong positions on protecting participant data by not sharing it with proprietary LLMs. P10 shared, “Obviously, I wouldn’t use my participant data with GPT-4 and a lot of these other models where I don’t know how they’re taking care of the data.” P18 expressed a similar strength of conviction: “I can’t feed my data into ChatGPT, or something like that, because of confidentiality concerns.” In the absence of clear, universal guidelines from IRBs within universities or venues providing recommendations for these issues, some researchers are creating their own strong norms: “I never put the data that is private to the model,” P5 confirmed. P7 and P13 highlighted that it is currently challenging to understand if LLM-based features are integrated into existing QDA tools in a way that protects participant data.
For example, P13 noted that MaxQDA now has LLM-based features, but their team chose not to use them because they were unable to confirm with their university whether using these tools was legal based on their data agreement and privacy standards. P7 echoed this, sharing their concern is not “just the researcher directly using ChatGPT, but also the researcher using tools which use ChatGPT further.” Researchers were also concerned these risks may be exacerbated for vulnerable populations, groups which are often the subject of study in qualitative research. Researchers in our study work with vulnerable populations, such as undocumented immigrants, those in political exile, and those with disabilities. Each one of these participants shared explicit concern that using LLMs on their participants’ data, given the populations they study, could put their participants risk, so they felt their data should not be used with LLMs.
To mitigate these risks, some researchers reported experimenting with ways to honor privacy while still using these tools. P11 shared, “Sometimes I do put some participant’s quotations in, but I intentionally trim out specific keywords, for example, specific product names or functions.” P9 said researchers should take proactive stance towards participant privacy: “Am I allowed to upload even anonymized content to an LLM? Usually you don’t sign a consent specifically on that question. And maybe legally it’s fine. But I feel like it shouldn’t be shared with a third party that might use it as a training tool.” Several proposed a stronger change to consent practices in anticipation of these concerns, like P7, who said “I think it’s of the utmost importance to inform the participants of the research and let them know we would be following this method or doing this type of analysis wherein we are sharing some of their information with the LLMs.” Broadly, in the absence of strong guidelines, researchers appeared to genuinely grapple with appropriate use of these tools while still following their own convictions and research norms.
4.2.4. Concerns around validity and evaluation
Participants considered whether LLMs would produce valid, usable outputs for research. Participants cited that LLMs produce hallucinations and incorrect information. “LLMs can give you wrong information” said P18, and P4 agreed: “some points are just wrong.” More broadly, participants often found that LLM, specifically ChatGPT, outputs were vague, odd, or missing what they considered key aspects: “they’re getting some of these broad ideas, but maybe the variables or the concepts that are in this theory are a little off or they might be different theories,” P6 shared, and P16 said “the LLMs tend to write really generically and I would have concerns.” They sometimes conclude these tools may not work for their qualitative research due to these limitations. Researchers shared that the practice of collecting data (like interviews) themselves gave them an intimate understanding of the data, which could help them guard against misinterpretation of the data through an LLM’s hallucination or misconstrual (P16, P20).
Researchers also reported uncertainty regarding the best way to validate and evaluate LLM performance in a way that aligns with their research community. P18 shared that LLMs “are all constantly being updated […] So the model is always changing, and that means that the output that you get is always changing.” In response to these concerns about validity and reproducibility, many participants emphasized that checking the LLM’s outputs is key being confident in the use of these tools in research (P5, P7, P10, P14, P18, P19, P20) and some others also mention that it is unacceptable to not check outputs (P1, P5, P18). P10, for example, emphasized the need for researchers to take responsibility: “there needs to be some sort of oversight because there will be an error or mistake through using LLMs, and it’s on the researcher to catch those errors or to have some safeguard.”
4.2.5. Concerns about model bias
Researchers raised concerns that biases embedded in LLMs might make them unsuitable for qualitative research on populations not well-represented in training data (P6, P7, P8, P10, P12, P16, P18, P19) including non-English-speaking populations (P1, P12, P18, P19). Are LLMs “able to understand data from someone with a particular identity?” P8 wondered. Researchers reported both experiences with and fears of LLMs misinterpreting data from marginalized groups through the mistranscription of important language from that community, or by making judgments that are not faithful to the community. For example, one participant noted “AI often falsely identifies queer content as hateful,” confirming that an LLM’s construals of their participant group may be at odds with faithful stewardship of that community’s interests. Another participant noted ChatGPT gave reductive information about women from different backgrounds experiencing grief when they asked ChatGPT. “Probably, if you’re saying, ‘what issues do women experience?’ [ChatGPT is] going to try and give you the most broad, all encompassing answer. There’s not really any way to be sure that what it’s giving you is covering all the ground, which is why I think it’s useful for supplement.” P19 cited concerns over known Western bias as the main reason they do not feel comfortable using LLMs in their work: “We all know that these models are pre-trained on datasets that are highly Western, mostly completely built from universities and companies that are based in developed places.” On their interest in using LLMs in qualitative research, they continued, “So for now, no, in the future I don’t know, maybe. We have a long way to go for non-English speaking languages.” Researchers worried that these LLM biases could perpetuate existing power inequalities or harms towards marginalized groups. P6 observed, “The machine is made within these power structures… so inherently it’s going to privilege some things over others because that’s how the world works,” a concern also shared by P12. Noticing similar properties of LLM training data, P7 said, “Most of the literature that’s accessible is, is the voice of the powerful, right?” Researchers were attuned to the connection between what data LLMs have access to and what they can faithfully represent, a property creating a potential mismatch between the properties of LLMs and researchers’ desire to faithfully analyze and represent their subject populations.
4.3. Qualitative Research and LLMs: Inherent Tensions?
Simultaneous with curiosity and exploration of LLM tools, participants grappled with potential tensions between the values of qualitative research and the capabilities LLMs may enable.
4.3.1. Emergent insights from the data, versus views imposed by LLMs
Many qualitative researchers use methods where findings emerge from the bottom up from the data, such as grounded theory, which encourages interactions with data without imposing a view from pre-conceived theories. Researchers wondered whether bringing in the external interpretations, language, and potentially even theory from an LLM were in alignment with qualitative approaches that value a researcher’s bottom up approach (P1, P3, P6, P9, P12, P16). P6 worried about letting LLMs lead theme generation: “My biggest worry is having AI code qualitative data: that feels a little bit different than engaging in a thematic analysis where we’re letting the data drive the themes and we’re recognizing that as humans, we’re bringing the meaning to those themes and putting that in.” Using LLMs to create an interpretation of data could also lead to the reproduction of standpoints in the LLM’s training data that marginalizes people through stance, or produces outright errors. P16 said, ”these people didn’t agree to editing their words or modifying their perspective” stressing the personal responsibility researchers feel in making sure participants are heard on their own terms. On the other hand, past tools have supported researchers in exploring data in new ways (Jiang et al., 2021) and builders could design tools to keep the values of exploration and emergence in mind.
Relying on interpretations produced by an LLM may also reduce researchers’ engagement with their own context and subjectivity. P1 shared: “There’s a little bit of like domain expertise that you’re able to apply as a researcher and spend time in the space and spend time with these individuals,” and in using LLMs to do this, P16 said, “I think we lose out on that, because now we’re going with this very centralized, generalized model.”
If LLMs present researchers with a single ”generalized” view of the data which they accept in their workflows at face value, we risk losing diverse ways of eliciting insight from the data. Previous work has surfaced similar concern from qualitative researchers, who express wanting to develop their own ideas before being influenced by external insights and theories from AI, (Overney et al., 2024; Lam et al., 2024); Hong et al. (2022) similarly notes participant concerns that “algorithms that output potentially immutable data summaries could bias knowledge generation.” The single view presented by LLMs in typical interfaces like ChatGPT also do so without presenting likely or unlikely alternatives, but future interfaces could make exploring other potential outputs easier, potentially allowing researchers to see their data in new ways. Given that interactions with AI can durably shape our beliefs, (Costello et al., 2024) researchers understandably wonder how LLMs may shape their insights, especially if they interact with LLMs before deeply engaging with the data themselves.
LLMs generally collapse context by obfuscating sources of information in produced text, making it hard for researchers to assess both the validity of the outputs as well as the source of influence from the LLM on their process. This obfuscation may be improved by retrieval-augmented generation (RAG) systems that increasingly cite sources when generating LLM outputs, but as of now, systems can contribute to context collapse that may undermine validity as well as researchers’ own confidence in their findings.
4.3.2. Close engagement with the data, versus LLMs that automate processing
Participants reported that directly collecting interviews and reviewing them manually enables sensemaking and synthesis (P4, P6, P8, P10, P16, P19, P20). P20 showcased this perspective, saying “I do think there’s something in looking at all the transcripts, having done all the interviews yourself, and having intuitions around what themes arise.” P10 also felt manual sensemaking was valuable, and tied this view to a potential tradeoff: “You also wonder: if an AI were to do it, would you lose some of that sensemaking that naturally happens when you’re individually doing it yourself? And could that lead to not a strong findings or not a strong discussion section?” Relatedly, some researchers worry that the ubiquity and ease of using LLMs in these related tasks may lessen the degree of deep engagement with the data through manual sensemaking by allowing researchers to rely on ”shortcuts.” P19 shared that if someone is “using [an LLM] as a shortcut for reading large-scale things [where] you can’t go through what’s generated, then I do have concerns.”
P13 saw the main risks of involving LLMs in data analysis as only resolvable if researchers still take care to collect and review data themselves: “I’m sure there’s lots of ways this could get messed up, like only looking for certain words or words that go together and missing on nuance or interesting ways other people are talking about the same concept. If you’d never read it yourself, or were not actively engaged in the development of, say, the words, or whatever you’re looking for, any pattern, creation or recognition… maybe something would get lost.” Researchers wrestled with whether deep knowledge of the data would be maintained if LLMs were used in the analysis process, echoing past concerns about technological involvement in qualitative research.
4.3.3. Relationship and sense of responsibility to participants, versus an emotionless machine.
Deep engagement with participant data, as well as an often ongoing relationship with research subjects and their stories, means that many qualitative researchers internalize a responsibility to do justice to participants in their research process and to act in their best interests. P18 expressed a fear about how LLMs could change this relationship: “If people found out that I was feeding data into an LLM, I think that can cause a breakdown of trust. Some of my participants are people I intend to continue working with in a more long term way.” When considering using LLMs for data analysis, P1 hesitated, saying “we’re not necessarily building empathy in the same way, for participants”, and P4 told us “it’s not doing the data justice, you know, by just having an LLM look at it.”
4.3.4. Subjectivity and multiplicity, versus one hegemonic way of seeing
Participants valued the context researchers bring to their work. Participants acknowledged interactions with LLMs may uncover new insights about their data, but they were concerned those interactions could result in some insights being lost. For example, qualitative researchers see value in noticing things beyond the surface of transcribed interview text (P4, P6, P7, P10, P14, P16, P20), and note that LLMs do not see or know what humans do. P9 shared, ”I think you do gain a lot of insights by not just the words that are typed: the way I hesitate on certain things like accents, body movement… there’s a tacit knowledge of the qualitative process that is not easily formalized and programmed.” As P7 put it, qualitative researchers “situate the text in a context,” not just analyze the “text as it is.” This led P7 to conclude that “in such cases, the LLMs might not be very helpful.”
As helpful as LLMs may be in discovering patterns or deriving new insights from the data, researchers note that LLMs may change what humans notice in their data, potentially missing critical insights. P19 noticed: “ There’s just something interesting that I see that it doesn’t give me, and the most important point of my analysis is to figure out the themes. If they’re figured out for me, how do I know that I haven’t missed something that is potentially interesting?” Echoing a quandary regarding what an LLM would surface in data, P14 said ”I would look for a different weighing of priorities.” If researchers rely on LLMs to analyze data, and LLMs shape what humans notice in the data at the expense of insights humans would otherwise bring, some kinds of insight may be increasingly lost. P3 also wondered if LLMs would focus on different insights than humans, who collaboratively notice and integrate findings together: “I think when you ask for an LLM to give you a specific set of instructions, you lose some of that nuance and some of that discovery. And so I definitely think certain types of information could be lost or just never analyzed.”
4.3.5. Who is in charge or creating meaning from human experience?
Many researchers in our participant group discussed how qualitative methods are well-suited to understanding the why behind human experience, (P1, P2, P6, P10, P13, P16, P20), particularly in contrast to what quantitative approaches can tell us. One theme is resounding: qualitative researchers want to keep humans in charge and center their perspectives. Summarizing this point, P8 shared, “The whole point of doing qualitative research is to engage with people, and if we just stop doing that, then I feel like it endangers qualitative research in some way.” P15 took the view using LLMs in qualitative research may still be compatible with this goal as long as humans are the ones ultimately determining meaning from an analysis: ”I think LLMs would be incredible for topical emergence. Only humans can really [tell] you what the themes are ultimately.” Can LLMs surface useful complementary insight while we keep humans in charge of what those features mean in the context of human lives? LLMs do not have lived experience, and do not “understand” the lived experiences of humans. In a set of methodologies concerned with making sense of the “texture of human lives” (P19), keeping the project of creating meaning as a human-centered project is critical.
5. Discussion
Key themes emerged from interviews that allow us to contextualize the current moment, and imagine how qualitative researchers may adopt LLMs moving forward while staying faithful to the responsibilities they have towards participants. The qualitative researchers we interviewed were largely open to or interested in the conscientious use of LLMs across the research pipeline. Previous work has evaluated specific systems for qualitative researchers that leverage LLMs or other AI (Gao et al., 2023a; Gebreegziabher et al., 2023; Overney et al., 2024), finding promising use cases for qualitative researchers. In contrast, we studied participants’ engagement with LLMs as they themselves had chosen, independent of any one tool or prescribed use case. In participant answers, we found optimism about many potential uses but also valid concerns about how to align these uses to their personal and disciplinary research values. We found this across participants at many stages of familiarization (Long et al., 2024), between arization (not familiar with LLM tools), familiarization (actively playing with LLM tools), and post-familiarization (deploying LLMs for specific chosen use cases).
A second major finding is that although researchers are increasingly surrounded by task-focused AI tools, their imagination of “AI” or “LLMs” is dominated by the chatbot interaction pattern exemplified by ChatGPT. Participants are comfortable with task-specific tools that use LLMs such as grammar check, speech-to-text transcription, and translation, and use them extensively with generally (but not universally) high levels of trust. Their perception of the suitability of LLMs for research, however, is largely focused on the single ChatGPT system and the applications easiest for it, such as ideation and text editing; even among this more computational group, only the most technologically savvy participants were currently leveraging LLMs to do more complex tasks such as rigorous qualitative coding.
ChatGPT and its competitors are specifically designed to emulate a helpful assistant. For participants, this design decision increasingly blurs the boundary between seeing LLMs as a “tool” or as a “collaborator.” In contrast to previous contested technological advances in QDA software tools, ChatGPT is uniquely compelling because it performs human-like interaction patterns such as following natural language instructions, and gives the appearance of recognizing nuance in text described in abstract terms, not just counting words or providing statistics. Prior research supports this finding: the chatbot is an example of a user interface metaphor (Neale and Carroll, 1997); specific systems drive perceptions of broader technologies with Galloway’s interface effect (Galloway, 2012); and brief interactions with systems produce misleading impressions of overall performance (Bowman, [n. d.]). As these effects are unlikely to be limited to qualitative researchers, this work has potential impact in many domains beyond HCI.
On top of surfacing researchers’ points of excitement for future uses of LLMs to assist their research, we outline the many concerns they have, including uncertainties about norms, privacy, and output quality that could diminish the research process or misrepresent participants. Worries about norms and guidelines are ubiquitous, despite the active development and publication of norms by institutions and publication venues (e.g. ACM policies), ethical and supervisory boards, and academic research papers (Watkins, 2023; Hosseini and Horbach, 2023). In the social norm change literature, scholars posit that there is often a tipping point where social norms change rapidly (Andrighetto and Vriens, 2022). Our research suggests that some norms regarding LLMs may be close to a clear tipping point, such as the understanding that identifiable participant data ought not be sent to a third-party tool that may train AI models on the data, while norms regarding some other concerns have so far eluded convergence. Our work confirms prior perspectives that AI evolution is outpacing our norm formation processes (Baronchelli, 2024; Feldstein, 2024), and we show how scholars feel pressure and uncertainty when justifying LLM-aided techniques for academic publications. We show these systems being adopted despite wide-ranging concerns about privacy, suggesting a structural failure to educate researchers on best practices—a concern likely implicating research across fields, agnostic to methodology.
Beyond these practical concerns, a set of fundamental tensions arose in considering the use of LLMs to conduct qualitative research. There has long been a tension between work seeking to combine computational power with interpretive approaches (Baumer et al., 2017; Nelson, 2020) and work arguing that qualitative values must be protected from quantitative evaluation standards (Crabtree, 2024; Soden et al., 2024). Our interviews suggest that LLMs may continue blurring the line between positivist and interpretive approaches, as exemplified in emergent work (Lam et al., 2024; Choksi et al., 2024). As Messeri and Crockett (2024) discuss, the way some approaches are more readily aided by LLMs may change the research landscape by making some questions more tractable than others. There is a danger that the ability to analyze data at scale with less human involvement may make situated, slow, contextual and interpretivist approaches to qualitative data less attractive because of quick publication cycles and the perceived objectivity of LLMs. Mixed method researchers, who make up roughly half of our participants, may be positioned to accelerate this change. Published work uses LLMs for many types of text analysis, including in powerful new tools that are increasingly positioned to facilitate “theory-driven analysis” of textual data (Lam et al., 2024) and they may be adopted by users regardless of methodology. As LLMs produce summaries, topics, and labels of textual data that look increasingly fluent, nuanced, and contextual, researchers may increasingly be tempted to use these outputs in analysis while decreasing the type of engagement with participants and their data that researchers say is central to qualitative research values.
6. Limitations
While we recruited participants across disciplines, our participant group does not cover all perspectives on qualitative research and its values. Disciplines that are even more situated than those represented in our participant group, such as ethnography, may be at greater risk of diminishment by LLM involvement, and those perspectives may not be represented by the voices highlighted here. We believe the more contextually situated, slower, and interpretivist qualitative epistemologies are, the more LLMs are likely to be at odds with or threaten their research values. We also recruited participants regardless of their level of engagement with LLMs as users or as researchers. This elicited important findings about current use by non-experts, important intuitions from qualitative researchers regardless of their level of use, and insights from experts on responsible AI who avoid LLMs. However, some of these researchers, by virtue of not being experts on LLMs and their potential flaws, may have presented a different set of concerns than experts, particularly on the nature of risks from LLMs or their potential capabilities under different conditions of use. The fact our participants were mostly PhD students also likely shapes the set of concerns elicited, level of engagement with these tools, and interest in future use of LLMs in research. We also note our participants largely did not surface experience or opinions on simulated research subjects or automated interviewing, which are garnering significant attention in other research circles.
7. Implications
In this section, we consider the practical implications of our findings for two audiences: LLM-curious qualitative researchers, who are wondering if and how qualitative data work can be conducted ethically and effectively, and future qualitative data analysis tool builders, who are contending with how to build tools that facilitate these experiences.
7.1. Guidelines for researchers using LLMs in qualitative work
Based on our interview findings and prior work, we outline a set of guidelines for qualitative researchers interested in using LLMs in their research process. Qualitative researchers are currently already using and likely expanding their use of LLMs in research. The goal of these guidelines is to scaffold an approach for conscientious use of these tools beyond the current moment for researchers who may be considering whether and how to use them.
7.1.1. Intentionally select tools
Most participants who had used LLMs referred to ChatGPT, which is the most popular LLM interface currently available. However, ChatGPT is not an interface designed to be used for qualitative research. Understanding what data a tool was trained on, whether it responds well to research-related prompts, and whether it handles data appropriately. While there are not currently bespoke LLM models for qualitative research, they might exist in the future, and dedicated analysis platforms are incorporating LLM-based features.
7.1.2. Follow venue and institutional policies
Several venues have created policies for appropriate use and disclosure of LLMs in research, including the ACM and *ACL conferences (noa, [n. d.]; Chairs, [n. d.]). For example, the ACM allows use of generative AI, but a manuscript cannot be “primarily the result of the tool’s generative capabilities,” and authors must disclose generative AI use in the paper. The publishing venue Sage has similar guidelines, allowing use for editing but banning reviewers from inappropriately using generative AI (Sage, 2024). Follow your community’s guidelines, and consider how you might shape them going forward.
Beyond adhering to institutional guidelines mentioned above, when using LLMs as part of an analysis method or for analysis, disclose use fully and be explicit about where models were used, including which model was used and any fine-tuning, optimization, or prompt engineering which was performed. You may wish to engage in additional disclosure, particularly for uses in qualitative research that may have shaped original contributions forming theory on the data, like the generation of themes from the data using an LLM. On the other hand, in the absence of explicit venue or university guidelines, you may decide that light editing is appropriate without disclosure, such as the use of Grammarly to lightly edit texts, if tools did not shape the theoretical contribution or framing of the work.
7.1.3. Validate outputs for your use case
Prioritize validation of LLM outputs in proportion to the significance of a result. Our results raise concerns that the core intellectual contribution of qualitative research may be subject to the temptation to trust an LLM “collaborator.” Regardless of methodology, you must remain responsible for your process and results. You should choose methods of validating the use of LLMs according to your use case and methodological values, selecting tools and methods that are transparent and disclosing your process as much as possible. When using LLMs to produce small-$N$ outputs, such as a single interview protocol or translation, you should ensure these are carefully checked and reviewed. If using LLMs to produce large-$N$ outputs, such as applying qualitative codes at scale or deploying personalized interview guides, you should have a plan to validate a subsample of the data and check performance on your task. Furthermore, you should consider whether more transparent methods than using LLMs for your chosen task exist, and if so, prioritize using them, since current lack of transparency and explainability of LLM outputs makes transparency in the research process more difficult. Whether or not LLM outputs can be validated according to an existing standard may depend on the chosen epistemology. For example, depending on your research values, you may decide not to use LLM-generated codes, if you decide that your research places value on human-generated perspectives on coding instead. You should also consider whether you may be committing accidental acts of academic plagiarism by relying on LLM-generated outputs (Bhardwaz and Kumar, 2023) across the research pipeline, and should take responsibility for checking outputs for any copied content accordingly.
7.1.4. Protect participant data
Qualitative researchers must protect participant data and anonymity. Across the board, but especially if using a closed source or proprietary model, take extreme care not to share personally identifiable information or compromise the data of at-risk populations. Consider whether hosting open source models locally, privately, and securely is feasible based on your project constraints as well as suitable for your use case. If you plan to use a proprietary model, carefully consider how you might reliably anonymize data and whether participants have been given the chance to provide explicit consent to have their data shared with third-party models for analysis. Avoid LLM interfaces which will or could use your participants’ data as training data, given risks to their interests, without their consent. Consider providing explicit justification if using a proprietary model for scientific purposes, as recommended by  Palmer et al. (2024). If using an external software tool, such as NVivo or Atlas.ti, that incorporates LLM-based functions, check whether the tool is using a proprietary model that may change your assessment of its appropriate use with your data, either according to university policy or IRB standards. As a matter of process, update your consent forms for participants to accurately reflect any intended use of LLMs in the research and implications for participants. This should include details on how participant privacy will be protected (through using anonymized data on local models, for example, in contrast to companies that may use participant data to train future models).
7.1.5. Consider a model’s training data when evaluating potential harms
The major LLMs available today, such as ChatGPT, Claude, or Gemini, have been trained on large swaths of data on the Internet (Kandpal et al., 2023), but commercial models largely lack transparency on their data sources (Dubey et al., 2024; Soldaini et al., 2024). They perform best when applied to data that is closer to their training data, as our participants noted, but may appear confident and — on the surface — plausible, despite a lack of transparency into their mechanisms. However, the details of qualitative research processes are typically not published on the Internet to preserve participant privacy, nor is there a large amount of qualitative research data in general, so it is unknown but unlikely that major models have been trained in ways that are well-fitted to qualitative research praxis or to specialized groups of participants for tasks like coding data. They also may perform especially poorly on data concerning non-Western participants, non-English speaking participants, or people from otherwise non-dominant backgrounds. If your research, like much qualitative research, focuses on these groups, using LLMs could misalign with the goals of your research.
7.1.6. Consider potential risks to your sensemaking process
When considering how to use LLMs, consider whether the tool is doing the data and your research questions justice. For example, consider whether you may be using an LLM as a way to spend less time with the data, whether you are introducing a limited or biased view on the data, and whether you could become overreliant on the tool. All researchers should consider whether using a tool that changes the sensemaking process diminishes or enhances it, particularly depending on one’s research values.
7.2. Design principles for qualitative research tools incorporating LLMs
We have identified the urgent need for transparent tools that center qualitative research use cases. Participants already use chat-based proprietary LLM interfaces like ChatGPT for qualitative research tasks, while dedicated analysis tools, like NVivo, do not currently serve comparable roles or have comparable features. Conversely, general-purpose chat-based LLM interfaces are not designed as qualitative research tools. As such, we outline some use cases that present opportunity areas for the development of tools for qualitative researchers that leverage LLMs.
Based on our findings, tool builders ought to emphasize 1) interactivity, 2) data exploration, and 3) qualitative coding in future qualitative research tools. Participants consistently mentioned the useful nature of the interactivity LLMs afford through a chat-based interface, a new modality compared to standard analysis software tools. Participants highlighted that chat interactions help them ideate, obtain feedback, and that they would appreciate making requests for data retrieval or formatting in natural language as is increasingly possible  (Izacard and Grave, 2021; Gao et al., 2023b). There is a gap in tools that facilitate these use cases seamlessly integrated with other desired uses. Second, researchers also perceive an opportunity to use LLMs to explore their data or see it in a new way. LLMs could help researchers see their data through a multiplicity of lenses, “support serendipity” (Jiang et al., 2021) for qualitative researchers and ideally widen insight rather than homogenize it. Lastly, qualitative researchers are interested in using LLMs to conscientiously assist in the coding process. Tools could also help scaffold the qualitative coding process for beginners, but some worry that tools may diminish ”essential friction” (P15) as one learns coding. Tools should help researchers check if the LLM is implementing their desired definitions and codes correctly. Across all opportunity areas, implementing LLM-based features also present risks which must be navigated carefully, such as the risks of anthropomorphizing AI through interactivity, becoming overreliant on it, decreasing researcher agency, reducing novelty, or the risk of broader misalignment with qualitative research values.
Researchers want to use tools that align with their values, and some implementations could empower researchers, while others may risk dispossessing them or participants. Below are a few principles for tool designers when considering how and if to create tools that responsibly address these opportunity areas:
7.2.1. Design for participant privacy
Proprietary LLMs have sparked warranted concerns regarding participant privacy. Some researchers mention closed source models as a potential solution, but most are unable to implement solutions to this problem themselves. Making it easy to host open-source models for some qualitative research functions would help qualitative researchers feel more confident about using LLMs in qualitative research. Tools that incorporate LLMs should be explicit about when and how they call external APIs, how they handle privacy, how they expect users to deanonymize data before using tools, how they scrub or plan to delete data, and whether they use inputs for training models.
7.2.2. Design for validation
Different epistemologies will have different ideas about what it means to validate outputs. Part of designing for researchers to feel agency and trust is giving them ways to validate whether the use of LLMs in a process gives useful and valid outputs and make evaluation transparent to let researchers feel empowered to make choices and feel confident in their decisions throughout the research process. This includes designing features to understand or probe outcomes and suggestions, as well as designing for reproducibility of outcomes when possible. Features that require researchers to review outputs depending on the task may help accomplish this goal.
7.2.3. Design for agency
Researchers want to maintain agency over if, when, and how they are assisted or influenced by AI tools in their research process. One way to design for agency is to give researchers ways of opting in or out of AI-based features within platforms, facilitating researchers’ own development of theory from the data before providing assistance. Designing for agency also involves designing against overreliance on or anthropomorphism of these technologies.
7.2.4. Design for researcher context
Researchers often wish to bring in their own context when performing qualitative analysis, including their past experiences, the texts that influence their work, and the theories in which they ground their work. Tools for qualitative researchers should be designed assuming researchers have unique perspective to contribute to the data.
7.2.5. Design for deep engagement with data
Qualitative researchers worry that LLM-based tools could distance them from data rather than deepen their relationship with it, but tools can be explicitly designed to elicit deeper engagement with the data. Many qualitative researchers expressed interest in using LLMs to see their data through another lens, but found that insights and themes produced on their data are generic or repetitive. Tools can explore LLMs as a way to produce more unique ways of understanding their data, drawing researchers’ attention to less predictable features of the data rather than predictable ones, or helping researchers examine or challenge their existing theories with direct evidence from the data, ideally deepening researchers’ understanding of the data rather than diminishing it.
7.2.6. Design to consider participant perspectives and interests
Researchers are concerned about the variable performance LLMs have across contexts, knowledge domains, cultures, and languages. Tools should work to facilitate better alignment between LLMs and diverse subject populations. When possible, tool builders could give the option to use models that better reflect a target, like models that specialize in a particular language or knowledge domain, and facilitate investigating where there might be errors or bias, in alignment with the recommendation to design for agency and validation.
8. Conclusion
Involving LLMs in qualitative research presents potential opportunities for augmenting research with new capabilities, but also surfaces tensions that show some types of LLM involvement may misalign with the values of qualitative research. Careful engagement with these problems as researchers, and thoughtful tool design from developers, can help navigate these tensions in order to both align with research values and support researchers in their intended use case. ChatGPT spurred new ideas for researchers regarding how their research processes might be augmented, but it should not serve as either a model or an endpoint for what is possible for LLMs in research tooling. Instead, we can integrate concerns, opportunities, and lessons to keep qualitative research values at the center of our work. The guidelines and design principles we suggest can help qualitative researchers and those in related fields, including journalism and many social sciences, center human perspectives to keep doing participants and their data justice as we navigate a new era of tooling.
9. Acknowledgements
We thank James Eschrich for their formative contributions to the scope, literature, and analysis in this piece. We also thank Mor Naaman, Emily Tseng, and Louise Barkhuus for their perspectives on this work. We thank the participants of the “LLMs as Research Tools” workshop at CHI 2024 for their contributions that informed the direction of this work.